# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v1

@@ C.2 Extended results on all models
     Llama-3.1-405B-Inst 0.159 $\pm$ 0.008 6.369 $\pm$ 0.384
     Llama-3-70B-Inst 0.518 $\pm$ 0.030 17.350 $\pm$ 0.399
     Llama-3-8B-Inst 1.647 $\pm$ 0.069 27.425 $\pm$ 0.360
     Mixtral-8x7B-Inst 1.061 $\pm$ 0.058 24.010 $\pm$ 0.499
     GPT-4o 0.276 $\pm$ 0.010 6.973 $\pm$ 0.304
     GPT-4o-mini 0.353 $\pm$ 0.022 13.704 $\pm$ 0.315
@@
     Llama-3-70B-Inst 0.550 $\pm$ 0.013 13.469 $\pm$ 0.399
     Llama-3-70B 0.237 $\pm$ 0.006 9.646 $\pm$ 0.542
     Llama-3-8B-Inst 0.484 $\pm$ 0.010 14.465 $\pm$ 0.356
     Llama-3-8B 0.313 $\pm$ 0.023 14.497 $\pm$ 0.713
     Mixtral-8x7B-Inst 0.264 $\pm$ 0.004 12.505 $\pm$ 0.482
     Mixtral-8x7B 0.262 $\pm$ 0.008 12.826 $\pm$ 0.381
@@
     Llama-3.1-405B-Inst 0.473 $\pm$ 0.005 25.194 $\pm$ 0.219
     Llama-3-70B-Inst 0.714 $\pm$ 0.035 28.141 $\pm$ 0.348
     Llama-3-8B-Inst 1.900 $\pm$ 0.059 34.027 $\pm$ 0.265
     Mixtral-8x7B-Inst 0.847 $\pm$ 0.045 26.253 $\pm$ 0.494
     GPT-4o 0.441 $\pm$ 0.008 23.450 $\pm$ 0.289
     GPT-4o-mini 0.423 $\pm$ 0.006 26.131 $\pm$ 0.199
@@
     Llama-3-70B-Inst 0.378 $\pm$ 0.004 19.742 $\pm$ 0.349
     Llama-3-70B 0.312 $\pm$ 0.006 16.413 $\pm$ 0.376
     Llama-3-8B-Inst 0.503 $\pm$ 0.009 23.125 $\pm$ 0.337
     Llama-3-8B 0.345 $\pm$ 0.003 18.819 $\pm$ 0.278
     Mixtral-8x7B-Inst 0.383 $\pm$ 0.015 18.454 $\pm$ 0.361
     Mixtral-8x7B 0.306 $\pm$ 0.007 17.097 $\pm$ 0.386
@@
     Lag-Llama 0.329 $\pm$ 0.004 22.781 $\pm$ 0.626
     Chronos-Tiny 0.328 $\pm$ 0.001 20.783 $\pm$ 0.335
     Chronos-Mini 0.341 $\pm$ 0.001 21.926 $\pm$ 0.331
     Chronos-Small 0.328 $\pm$ 0.002 20.126 $\pm$ 0.269
     Chronos-Base 0.672 $\pm$ 0.003 22.519 $\pm$ 0.283
     Chronos-Large 0.326 $\pm$ 0.002 19.255 $\pm$ 0.311
     TimeGEN 0.354 $\pm$ 0.000 26.700 $\pm$ 0.121
     Moirai-Small 0.565 $\pm$ 0.031 26.323 $\pm$ 0.353
     Moirai-Base 0.624 $\pm$ 0.013 26.065 $\pm$ 0.275
     Moirai-Large 0.520 $\pm$ 0.006 21.041 $\pm$ 0.718

### v2

@@ C.2 Full results partitioned by types of context
    Llama-3.1-405B-Inst 0.159 $\pm$ 0.008 7.455 $\pm$ 0.585 0.174 $\pm$ 0.010 0.146 $\pm$ 0.001 0.066 $\pm$ 0.004 0.168 $\pm$ 0.010 0.398 $\pm$ 0.045
    Llama-3-70B-Inst 0.286 $\pm$ 0.004 14.082 $\pm$ 0.194 0.336 $\pm$ 0.006 0.180 $\pm$ 0.003 0.185 $\pm$ 0.006 0.230 $\pm$ 0.005 0.629 $\pm$ 0.019
    Llama-3-8B-Inst 0.461 $\pm$ 0.008 26.903 $\pm$ 0.558 0.571 $\pm$ 0.011 0.313 $\pm$ 0.008 0.241 $\pm$ 0.015 0.263 $\pm$ 0.003 0.531 $\pm$ 0.005
    Mixtral-8x7B-Inst 0.523 $\pm$ 0.023 31.390 $\pm$ 0.393 0.724 $\pm$ 0.037 0.236 $\pm$ 0.002 0.185 $\pm$ 0.006 0.384 $\pm$ 0.028 0.875 $\pm$ 0.128
    Qwen-2.5-7B-Inst 0.289 $\pm$ 0.003 21.792 $\pm$ 0.725 0.289 $\pm$ 0.004 0.176 $\pm$ 0.003 0.281 $\pm$ 0.007 0.238 $\pm$ 0.002 0.525 $\pm$ 0.003
    Qwen-2.5-1.5B-Inst 0.575 $\pm$ 0.014 30.634 $\pm$ 0.834 0.682 $\pm$ 0.023 0.284 $\pm$ 0.006 0.358 $\pm$ 0.009 0.453 $\pm$ 0.004 1.270 $\pm$ 0.009
    Qwen-2.5-0.5B-Inst 0.464 $\pm$ 0.012 26.254 $\pm$ 0.514 0.610 $\pm$ 0.019 0.165 $\pm$ 0.004 0.221 $\pm$ 0.011 0.483 $\pm$ 0.016 0.429 $\pm$ 0.006
    GPT-4o 0.276 $\pm$ 0.010 8.609 $\pm$ 0.402 0.220 $\pm$ 0.007 0.118 $\pm$ 0.001 0.113 $\pm$ 0.001 0.258 $\pm$ 0.012 0.858 $\pm$ 0.053
    GPT-4o-mini 0.353 $\pm$ 0.022 16.463 $\pm$ 0.463 0.474 $\pm$ 0.035 0.139 $\pm$ 0.002 0.139 $\pm$ 0.002 0.343 $\pm$ 0.029 0.644 $\pm$ 0.128
@@
    Llama-3-70B-Inst 0.550 $\pm$ 0.013 17.042 $\pm$ 0.533 0.455 $\pm$ 0.018 0.516 $\pm$ 0.028 0.751 $\pm$ 0.022 0.573 $\pm$ 0.017 0.392 $\pm$ 0.028
    Llama-3-70B 0.237 $\pm$ 0.006 11.962 $\pm$ 0.676 0.213 $\pm$ 0.005 0.121 $\pm$ 0.008 0.282 $\pm$ 0.016 0.194 $\pm$ 0.004 0.360 $\pm$ 0.011
    Llama-3-8B-Inst 0.484 $\pm$ 0.010 18.216 $\pm$ 0.449 0.477 $\pm$ 0.013 0.161 $\pm$ 0.006 0.302 $\pm$ 0.003 0.311 $\pm$ 0.008 0.878 $\pm$ 0.035
    Llama-3-8B 0.313 $\pm$ 0.023 18.373 $\pm$ 0.949 0.334 $\pm$ 0.035 0.123 $\pm$ 0.004 0.269 $\pm$ 0.016 0.287 $\pm$ 0.030 0.739 $\pm$ 0.134
    Mixtral-8x7B-Inst 0.264 $\pm$ 0.004 15.207 $\pm$ 0.611 0.242 $\pm$ 0.007 0.173 $\pm$ 0.004 0.312 $\pm$ 0.012 0.217 $\pm$ 0.002 0.437 $\pm$ 0.007
    Mixtral-8x7B 0.262 $\pm$ 0.008 15.902 $\pm$ 0.488 0.250 $\pm$ 0.008 0.119 $\pm$ 0.003 0.297 $\pm$ 0.018 0.227 $\pm$ 0.007 0.457 $\pm$ 0.011
    Qwen-2.5-7B-Inst 1.980 $\pm$ 0.027 42.510 $\pm$ 0.710 2.518 $\pm$ 0.044 2.857 $\pm$ 0.056 1.891 $\pm$ 0.008 1.605 $\pm$ 0.036 1.333 $\pm$ 0.144
    Qwen-2.5-7B 0.904 $\pm$ 0.037 35.498 $\pm$ 0.982 1.140 $\pm$ 0.047 1.002 $\pm$ 0.053 0.594 $\pm$ 0.068 0.633 $\pm$ 0.048 0.928 $\pm$ 0.129
    Qwen-2.5-1.5B-Inst 2.153 $\pm$ 0.027 47.557 $\pm$ 0.844 2.605 $\pm$ 0.041 1.672 $\pm$ 0.055 1.469 $\pm$ 0.023 2.054 $\pm$ 0.034 2.448 $\pm$ 0.128
    Qwen-2.5-1.5B 1.731 $\pm$ 0.036 42.332 $\pm$ 0.511 2.337 $\pm$ 0.049 2.982 $\pm$ 0.052 0.876 $\pm$ 0.067 1.489 $\pm$ 0.046 1.304 $\pm$ 0.129
    Qwen-2.5-0.5B-Inst 1.938 $\pm$ 0.024 47.511 $\pm$ 0.580 2.445 $\pm$ 0.038 1.960 $\pm$ 0.063 1.563 $\pm$ 0.009 1.762 $\pm$ 0.031 1.199 $\pm$ 0.129
    Qwen-2.5-0.5B 1.991 $\pm$ 0.024 45.725 $\pm$ 0.798 2.539 $\pm$ 0.039 2.083 $\pm$ 0.052 1.651 $\pm$ 0.013 1.796 $\pm$ 0.031 1.225 $\pm$ 0.128
@@
    Llama-3.1-405B-Inst 0.473 $\pm$ 0.005 32.451 $\pm$ 0.281 0.527 $\pm$ 0.007 0.713 $\pm$ 0.014 0.375 $\pm$ 0.003 0.321 $\pm$ 0.002 0.587 $\pm$ 0.005
    Llama-3-70B-Inst 0.470 $\pm$ 0.008 33.490 $\pm$ 0.345 0.532 $\pm$ 0.013 0.676 $\pm$ 0.018 0.370 $\pm$ 0.003 0.316 $\pm$ 0.002 0.615 $\pm$ 0.005
    Llama-3-8B-Inst 0.601 $\pm$ 0.006 39.660 $\pm$ 0.383 0.747 $\pm$ 0.009 0.679 $\pm$ 0.015 0.330 $\pm$ 0.008 0.333 $\pm$ 0.003 0.604 $\pm$ 0.004
    Mixtral-8x7B-Inst 0.712 $\pm$ 0.021 44.243 $\pm$ 0.349 0.906 $\pm$ 0.035 0.758 $\pm$ 0.015 0.382 $\pm$ 0.002 0.487 $\pm$ 0.028 0.893 $\pm$ 0.128
    Qwen-2.5-7B-Inst 0.420 $\pm$ 0.022 35.112 $\pm$ 0.521 0.478 $\pm$ 0.035 0.515 $\pm$ 0.017 0.316 $\pm$ 0.008 0.358 $\pm$ 0.029 0.830 $\pm$ 0.128
    Qwen-2.5-1.5B-Inst 0.450 $\pm$ 0.007 34.456 $\pm$ 0.591 0.493 $\pm$ 0.011 0.324 $\pm$ 0.007 0.365 $\pm$ 0.008 0.312 $\pm$ 0.006 0.498 $\pm$ 0.009
    Qwen-2.5-0.5B-Inst 0.305 $\pm$ 0.007 25.477 $\pm$ 0.327 0.341 $\pm$ 0.007 0.185 $\pm$ 0.004 0.237 $\pm$ 0.015 0.254 $\pm$ 0.005 0.396 $\pm$ 0.002
    GPT-4o 0.441 $\pm$ 0.008 29.851 $\pm$ 0.362 0.491 $\pm$ 0.013 0.280 $\pm$ 0.007 0.360 $\pm$ 0.002 0.274 $\pm$ 0.001 0.504 $\pm$ 0.002
    GPT-4o-mini 0.423 $\pm$ 0.006 34.043 $\pm$ 0.264 0.481 $\pm$ 0.009 0.391 $\pm$ 0.007 0.326 $\pm$ 0.004 0.280 $\pm$ 0.001 0.531 $\pm$ 0.003
@@
    Llama-3-70B-Inst 0.378 $\pm$ 0.004 24.668 $\pm$ 0.442 0.405 $\pm$ 0.006 0.186 $\pm$ 0.004 0.336 $\pm$ 0.003 0.251 $\pm$ 0.002 0.481 $\pm$ 0.004
    Llama-3-70B 0.312 $\pm$ 0.006 20.968 $\pm$ 0.437 0.312 $\pm$ 0.004 0.142 $\pm$ 0.004 0.306 $\pm$ 0.016 0.245 $\pm$ 0.002 0.479 $\pm$ 0.006
    Llama-3-8B-Inst 0.503 $\pm$ 0.009 29.622 $\pm$ 0.417 0.598 $\pm$ 0.014 0.262 $\pm$ 0.009 0.348 $\pm$ 0.004 0.265 $\pm$ 0.002 0.510 $\pm$ 0.001
    Llama-3-8B 0.345 $\pm$ 0.003 23.980 $\pm$ 0.358 0.387 $\pm$ 0.004 0.162 $\pm$ 0.006 0.260 $\pm$ 0.006 0.249 $\pm$ 0.001 0.491 $\pm$ 0.002
    Mixtral-8x7B-Inst 0.383 $\pm$ 0.015 23.531 $\pm$ 0.434 0.420 $\pm$ 0.024 0.162 $\pm$ 0.008 0.325 $\pm$ 0.004 0.351 $\pm$ 0.020 0.470 $\pm$ 0.005
    Mixtral-8x7B 0.306 $\pm$ 0.007 21.671 $\pm$ 0.467 0.295 $\pm$ 0.004 0.150 $\pm$ 0.004 0.320 $\pm$ 0.019 0.241 $\pm$ 0.001 0.489 $\pm$ 0.003
    Qwen-2.5-7B-Inst 1.023 $\pm$ 0.026 37.614 $\pm$ 1.153 1.440 $\pm$ 0.041 0.889 $\pm$ 0.032 0.362 $\pm$ 0.017 0.833 $\pm$ 0.033 0.810 $\pm$ 0.128
    Qwen-2.5-7B 0.732 $\pm$ 0.030 35.811 $\pm$ 1.119 0.924 $\pm$ 0.045 0.403 $\pm$ 0.034 0.426 $\pm$ 0.028 0.547 $\pm$ 0.035 0.792 $\pm$ 0.128
    Qwen-2.5-1.5B-Inst 1.515 $\pm$ 0.033 44.187 $\pm$ 1.036 2.107 $\pm$ 0.047 0.607 $\pm$ 0.038 1.113 $\pm$ 0.047 1.241 $\pm$ 0.042 0.926 $\pm$ 0.128
    Qwen-2.5-1.5B 1.069 $\pm$ 0.028 38.512 $\pm$ 1.027 1.296 $\pm$ 0.044 0.272 $\pm$ 0.019 0.779 $\pm$ 0.025 0.799 $\pm$ 0.035 0.785 $\pm$ 0.128
    Qwen-2.5-0.5B-Inst 1.318 $\pm$ 0.037 41.710 $\pm$ 0.714 1.536 $\pm$ 0.048 0.303 $\pm$ 0.014 1.224 $\pm$ 0.077 1.096 $\pm$ 0.044 0.803 $\pm$ 0.128
    Qwen-2.5-0.5B 1.819 $\pm$ 0.027 45.094 $\pm$ 0.692 2.250 $\pm$ 0.042 1.480 $\pm$ 0.054 1.603 $\pm$ 0.022 1.592 $\pm$ 0.035 1.004 $\pm$ 0.129
@@
    Lag-Llama 0.329 $\pm$ 0.004 29.076 $\pm$ 0.793 0.333 $\pm$ 0.005 0.167 $\pm$ 0.005 0.281 $\pm$ 0.008 0.299 $\pm$ 0.004 0.495 $\pm$ 0.014
    Chronos-Tiny 0.328 $\pm$ 0.001 26.277 $\pm$ 0.411 0.302 $\pm$ 0.002 0.163 $\pm$ 0.002 0.375 $\pm$ 0.002 0.263 $\pm$ 0.002 0.486 $\pm$ 0.003
    Chronos-Mini 0.341 $\pm$ 0.001 27.557 $\pm$ 0.392 0.318 $\pm$ 0.002 0.171 $\pm$ 0.003 0.389 $\pm$ 0.002 0.271 $\pm$ 0.002 0.481 $\pm$ 0.004
    Chronos-Small 0.328 $\pm$ 0.002 25.208 $\pm$ 0.350 0.308 $\pm$ 0.002 0.179 $\pm$ 0.002 0.375 $\pm$ 0.003 0.256 $\pm$ 0.002 0.453 $\pm$ 0.007
    Chronos-Base 0.672 $\pm$ 0.003 29.209 $\pm$ 0.361 0.570 $\pm$ 0.002 0.211 $\pm$ 0.005 0.372 $\pm$ 0.002 0.710 $\pm$ 0.003 2.481 $\pm$ 0.013
    Chronos-Large 0.326 $\pm$ 0.002 24.459 $\pm$ 0.381 0.314 $\pm$ 0.002 0.179 $\pm$ 0.003 0.360 $\pm$ 0.003 0.253 $\pm$ 0.002 0.460 $\pm$ 0.004
    TimeGEN 0.354 $\pm$ 0.000 34.434 $\pm$ 0.161 0.333 $\pm$ 0.000 0.177 $\pm$ 0.000 0.385 $\pm$ 0.000 0.292 $\pm$ 0.000 0.474 $\pm$ 0.000
    Moirai-small 0.565 $\pm$ 0.031 33.881 $\pm$ 0.420 0.662 $\pm$ 0.050 0.195 $\pm$ 0.010 0.418 $\pm$ 0.002 0.565 $\pm$ 0.041 0.464 $\pm$ 0.013
    Moirai-base 0.624 $\pm$ 0.013 33.146 $\pm$ 0.369 0.629 $\pm$ 0.021 0.172 $\pm$ 0.002 0.384 $\pm$ 0.004 0.640 $\pm$ 0.018 0.486 $\pm$ 0.015
    Moirai-large 0.520 $\pm$ 0.006 27.005 $\pm$ 0.857 0.596 $\pm$ 0.009 0.140 $\pm$ 0.001 0.414 $\pm$ 0.002 0.505 $\pm$ 0.007 0.438 $\pm$ 0.011